OpenStudy (anonymous):

Find the radius of a sphere if V = 330 and pi = 3.14. Answer 4.29 8.87 110 5.67

OpenStudy (anonymous):

good job you are right

OpenStudy (anonymous):

huh what do you mean i have not answered it yet

OpenStudy (anonymous):

23733.21951

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

V = (4 pi a^3)/3 V=330 (4 pi a^3)/3=330 (4 pi a^3)=330*3 a^3=(330*3)/(4*3.14)=78.82 a=4.29